OpenStudy (narissa):

How do you simplify 3√x^8?

OpenStudy (narissa):

@jhonyy9

OpenStudy (welshfella):

√x^8 = (x^8)^(1/2)

OpenStudy (welshfella):

use the rule (x^a)^b = x^ab

OpenStudy (narissa):

x^8(1/2) = 4? @welshfella

OpenStudy (anonymous):

not 4 will be x^4 but there in face of squart there is a 3 again - yes ?

OpenStudy (anonymous):

ok ?

OpenStudy (narissa):

x^8(1/2)=x^4? not 4 ok

OpenStudy (anonymous):

why you think that is just 4 ?

OpenStudy (welshfella):

yes also dont forget the 3

OpenStudy (narissa):

Because i added the numbers and got 4 i guess i didn't think to put x^4.

OpenStudy (narissa):

so now do we have 3√x^4?

OpenStudy (anonymous):

without sqrt

OpenStudy (anonymous):

sqrt was the exponent of 1/2

OpenStudy (anonymous):

bc sqrt(x) = x^(1/2) yes ?

OpenStudy (narissa):

yes

OpenStudy (anonymous):

so than there is sqrt(x^8) = (x^8)^(1/2) but you know that when there are exponent of exponent so than the exponents multiplie so will get (x)^(8*(1/2)) = (x)^(8/2) = x^4 ok. ?

OpenStudy (narissa):

Okay

OpenStudy (anonymous):

was my pleasure hope you understand it now but dont gorget the 3 in front of x

OpenStudy (anonymous):

forget - sorry

OpenStudy (narissa):

okay so my answer is 3√x^4? right?

OpenStudy (anonymous):

without sqrt just 3x^4

OpenStudy (narissa):

ohh okay thank you :)

OpenStudy (anonymous):

np was my pleasure bye bye